# Use as Many Surrogates as You Want: Selective Ensemble Attack to Unleash Transferability without Sacrificing Resource Efficiency

### v1

@@ 1 Introduction
Deep Neural Networks (DNNs) have been shown to be strikingly vulnerable to adversarial examples, i.e., maliciously crafted samples by adding small perturbations to benign inputs [35, 10]. An intriguing property of adversarial examples that makes them threatening in the real world is their transferability [1, 22, 45, 31, 16, 19, 47, 21], meaning that samples crafted on (known) surrogate models can also mislead (unknown) target models. Transferable attacks have been extensively studied, based on different strategies, such as gradient optimization, input transformation, and surrogate ensemble [49, 11].
Among them, surrogate ensemble is widely applied due to its simplicity and easy integration with other transfer techniques [40, 42, 50, 52]. In particular, there exists a well-known trade-off between attack transferability and the number of surrogate models, since using more models means lower resource efficiency regarding computational time and memory. Resource efficiency is even more critical when using transferable adversarial attacks for adversarial training-based defenses [38, 46, 34].
To avoid high resource overhead, the common practice is to simply restrict the number of models to only a few, typically $<5$ [6, 25, 30]. Under such a restricted condition, although complex, optimized ensemble strategies [44, 2, 37, 3] can somehow help, the resulting transferability is still far from using many more models, according to our experimental results in Section 4.2 and a concurrent study [27] on scaling law of transferability. In practice, it is meaningless to restrict the number of models since various pre-trained models are easily accessible online.
@@
In this paper, rather than proposing a new ensemble strategy, we argue that the trade-off between transferability and resource efficiency can be removed, i.e., we can adopt as many surrogate models as we want to improve transferability without sacrificing resource efficiency. This is made possible by our new interpretation of model diversity: decoupling model diversity into within-iteration and cross-iteration dimensions. In general, both dimensions account for transferability, but only within-iteration diversity (corresponding to the number of models used in each iteration) accounts for resource efficiency. Specifically, we quantify these two dimensions of diversity based on gradient similarity [17, 20] to explore how they respectively contribute to the final transferability.
@@
In sum, the main contributions of this work are as follows:
@@ 2 Related Work
Model ensemble attacks. Model ensemble attacks generally combine outputs from multiple models. Specifically, Liu et al. [28] averages loss from multiple models.  Dong et al. [6] further proposes logits or predicted probabilities ensemble. Advanced ensemble strategies are also studied. Specifically, Xiong et al. [44] notices differences between ensemble models and proposes Stochastic Variance Reduction Ensemble (SVRE) attacks to stabilize the gradient update.  Chen et al. [2] proposes an adaptive ensemble attack method (AdaEA), which adapts the models according to their specific contributions.  Tang et al. [37] introduces a novel method of Stochastic Mini-batch black-box attack with Ensemble Reweighing (SMER) to boost the transferability.  Chen et al. [3] uncovers two common weaknesses of model ensembles and relies on them to design the common Weakness Attacks (CWA). Although model ensemble attacks notably boost adversarial transferability, their practical utility is constrained by the substantial computational resources and time required for simultaneous gradient calculations across multiple models.
Gradient optimization transfer techniques. MI [6] introduces momentum into adversarial attacks, stabilizing gradient update directions. NI [25] further integrates Nesterov accelerated gradient into transfer-based attacks to better handle the history and future gradients. VT [39] utilizes variance tuning to find a more stable direction for gradient updates. PGN [9] relied on penalizing gradient norm to achieve flat local maxima.
Input transformation transfer techniques. DI [43] applies random resizing and padding operations to the input during each attack iteration to increase sample diversity and mitigate overfitting. TI [7] optimizes adversarial perturbations using a set of translated images and also reduces computational costs by convolving gradients of untranslated images with a kernel matrix. SI [25] follows a scale-invariant strategy, which optimizes the adversarial perturbations over the scale copies of the input image. SSA [30] transforms input images into the frequency domain and uses multiple input copies with diverse frequency features.
In this work, following the common practice in model ensemble attacks [44, 3], we consider integrating the above gradient optimization and input transformation transfer techniques. Note that another type of transfer technique, model refinement [42, 49], is excluded since it focuses on model modification rather than model ensemble.
@@ 3.1 Preliminaries of Model Ensemble Attacks
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,J(f({{x}^{adv}}),y),s.t.||{{x}^{adv}}-x|{{|}_{\infty}}\leq\varepsilon,$ (1)
where $\varepsilon$ denotes the constraint on adversarial perturbations using the $L_{\infty}$ norm. Instead of directly solving the complex optimization problem of Equation 1, the Fast Gradient Sign Method (FGSM) [10] and its iterative version, I-FGSM [18], adopt the projected gradient descent.
The model ensemble is commonly used to improve adversarial transferability based on the assumption that adversarial directions learned from a diverse set of surrogate models can better generalize/transfer to unknown target models. Early studies on model ensemble attacks [28] propose to average the loss values of $J(f({{x}^{adv}}),y)$ in Equation 2, resulting in the ensemble loss function as:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{n}{\frac{1}{n}J({{f}_{i}}({{x}^{adv}}),y)},$ (2)
where $n$ denotes the number of ensemble models. Differently, logit output values can be averaged:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{n}{J(\frac{1}{n}{{f}_{i}}({{x}^{adv}}),y)}.$ (3)
In this paper, we adopt the logit ensemble variant due to its simplicity and effectiveness [6].
@@ 3.2 Decoupling Within-iteration and Cross-iteration Model Diversity
Model ensemble attacks are known to yield improved transferability but require significant time and memory consumption due to the gradient computations on multiple models. Consequently, existing work has commonly restricted the number of models (typically $<5$) for an acceptable resource efficiency [9, 2]. To fully investigate the impact of the number of models on the attack transferability, we conduct exploratory experiments by using 20 different models (see detailed descriptions in Section 4.1). As can be seen from the experimental results in Figure 2, using more models consistently improves the transferability, despite the performance saturation when too many models are adopted. This finding is supported by [15], which verifies that increasing the number of models in adversarial attacks can reduce the upper bound of generalization error in empirical risk minimization.
@@
Figure 2: Attack success rates (%) of MI transferring from an ensemble of $m$ models to eight different target models. The result for TI in the Appendix E shows similar patterns.
Definition 1 Cross-iteration and Within-iteration Model Diversity. Given an iterative attack, the model diversity achieved by the number of distinct models used throughout the entire process, $n$, is termed cross-iteration model diversity, denoted by $D_{c}$. The model diversity achieved by the number of distinct models used per iteration, $m$, is termed within-iteration model diversity, denoted by $D_{i}$. Here $m$ remains a constant in each iteration and $m\leq n$.
Quantifying diversity. To explore the relation between the two dimensions of diversity in Definition 1 and actual transferability, we further quantify the diversity. Inspired by existing work on transferable attacks/defenses, we quantify them based on gradient similarity [17, 20]. Intuitively, different models yield different gradient directions on the input image. In an ensemble attack, collective gradient directions from different models enable adversarial examples to better traverse the decision boundaries of multiple models, thereby improving the possibility of transferring them to unknown target models. For within-iteration diversity, we compute the cosine similarity $S$ between gradients from different models in the same iteration:
$D_{i}=\frac{2}{m(m-1)}\sum_{1\leq i<j\leq m}S(\nabla_{x}J(f_{i}(x)),\nabla_{x}J(f_{j}(x)))$ (4)
@@
$D_{c}=S(\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t}}))}{m}},\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t+1}}))}{m}})$ (5)
@@ 3.3 Selective Ensemble Attack (SEA)
Algorithm 1 Selective Ensemble Attack (SEA) Algorithm
@@
1:  $\alpha=\varepsilon/T$; $x=x_{0}^{adv}$; $g_{0}=0$;
2:  for $t=0$ to $T-1$ do
3:     Randomly select $m$ models from set $\mathcal{F}$:${\mathcal{F}}_{m}$ = $\{f_{1},f_{2},...,f_{m}\}$ = $random.sample(\mathcal{F}_{s},m)$;
4:     Input $x$ and output ${{l}_{i}}(x_{t}^{adv})$ for $i=1,2,...,m$;
@@
Specifically, given $s$ easily accessible pre-trained models, denoted as $\mathcal{F}_{s}$ = {$f_{1}$, $f_{2}$, $f_{i}$, …, $f_{s}$}, SEA($s$,$m$) dynamically selects $m$ models in each iteration according to the required resource constraint. A larger $s$ generally leads to a larger number of cross-iteration models $n$, but $n$ may be smaller than $s$ when the iteration number is constrained (see detailed analyses of the expectation of $n$ in Appendix B). In contrast, the conventional method Ens($s$,$m$) has to limit $n$ to $m$. For model selection, we adopt the simple random selection and compare different sampling strategies in Table 9 of Section 4.3. The implementation details of SEA are presented in Algorithm 1.
@@ 4.1 Experimental Setups
Dataset and models. The experiments are conducted on the ImageNet-compatible dataset [32], which is widely used for evaluating adversarial examples [6, 25]. This dataset consists of 1000 images from 1000 ImageNet categories (one image per category), which can be correctly classified by all the adopted models described as follows.
For the target (test) model, we consider 2 CNNs, 2 ViTs, and 4 Defenses. Specifically, 2 CNNs are Inceptionv4 [36] and Resnet-v2-152 [12]. 2 ViTs are ViT-base [8] and Swin-base [29]. 4 Defenses are ens4-adv-Inception-v3 (ATens4), ens-adv-Inception-ResNet-v2 (ATens) [38], HGD [24], and RS [5]. For the surrogate models, we select them from 20 pre-trained PyTorch models with 8 diverse architectures: resnet34/50/101/152 [12], vgg11/13/16/19 [33], densenet121/161/169/201 [14], fbresnet152 [12], dpn98/107/131 [4], nasnetamobile [53], senet154, and se_resnet101/152 [13]. Notably, the setting of 4 surrogate models is consistent with that of previous works [25, 39].
@@
Table 3: Attack success rates (%) of Ens vs. SEA with eight different transfer baselines. The upper bound using all 20 accessible models in each iteration is denoted as Ens(20,20), which costs ×$\times 5$ more computational time and memory than Ens(20,4)/SEA(20,4).
@@
Transfer baselines and hyper-parameter settings. We integrate conventional model ensemble attacks (Ens) or our SEA with eight popular transfer techniques that are based on gradient optimization (i.e., MI [6], NI [25], VT [39], and PGN [9]) or input transformation (i.e., DI [43], TI [7], SI [25], and SSA [30]). We also consider more complex ensemble strategies, such as SVRE [44], AdaEA [2], SMER [37], and CWA [3]. All the above transfer baselines have been introduced in Section 2.
@@
Baseline Attack CNNs ViTs Defenses Avg. Time (min) Memory (GB)
@@
For clarity, we denote SEA and Ens under the ($s$,$m$) scenario as SEA($s$,$m$) and Ens($s$,$m$), respectively, where $s$ is the number of all easily accessible models and $m$ is the number of within-iteration models. When disregarding the architectural differences of models, the resource efficiency is directly proportional to $m$. For example, SEA(4,3) and Ens(4,3) are regarded as costing the same resource, and SEA(4,3) costs 1.5×1.5$\times$1.5 × resource of SEA(4,2). We also report the actual time and memory costs in Tables 4, 7, 8, 9 and in Appendix E, F, G. Additionally, the analysis of model loading time for SEA and Ens is reported in Appendix D.
Table 5: Attack success rates (%) of Ens vs. SEA on Google and Baidu Cloud Vision APIs with 100 randomly selected images.
Attack Google Baidu
Ens(4,3) 41 47
SEA(4,3) 44 51
Table 6: Attack success rates (%) of Ens vs. SEA on two commercial large vision-language models with 100 randomly selected images.
Attack Qwen ChatGLM
Ens(20,4) 40 39
SEA(20,4) 52 50
@@
Attack Surrogate models Target models Time (min) Memory (GB)
IncV3 IncV4 IncResV2 Res152 Avg. ATadv ATens3 ATens4 ATens Avg.
Ens(4,3) 86.4 86.3 78.6 79.2 82.6 13.0 12.9 34.3 28.3 22.1 38.5 18.3
SEA(4,3) 99.8 99.1 98.1 99.5 99.1 17.9 19.3 51.5 43.5 33.1 37.9 18.5
SEA(4,2) 100.0 99.0 98.2 99.5 99.2 16.9 17.9 47.3 39.5 30.4 25.7 16.1
Table 8: Transferability and resource efficiency of Ens vs. SEA with similar or diverse surrogates. In the similar cases, model ensemble is resnet18/34/50/101, vgg11/13/16/19, densenet121/161/169/201, or dpn68/98/107/131. In the diverse case, model ensemble is resnet101/vgg11/densenet161/dpn131. MI is the transfer baseline. Results for TI in the Appendix E show similar patterns.
Baseline Attack CNNs ViTs Defenses Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 94.8 98.2 51.2 51.3 74.5 63.5 90.9 62.0 73.3 1.8 7.3
Ens(4,3) 91.8 97.6 46.4 42.6 68.6 55.3 84.7 58.1 68.1 1.5 6.9
resnet SEA(4,3) 95.1 98.1 51.3 49.2 74.6 63.0 90.5 61.3 72.9 1.5 6.9
Ens(4,4) 99.9 93.5 45.4 43.8 64.2 50.9 74.5 52.3 65.6 2.6 10.6
Ens(4,3) 99.9 90.9 40.9 38.6 57.7 43.9 67.2 50.4 61.2 2.0 10.1
vgg SEA(4,3) 99.9 93.4 44.6 42.2 63.3 48.9 74.8 52.1 64.9 2.1 9.8
Ens(4,4) 92.7 99.9 58.3 56.3 78.6 71.0 91.7 63.2 76.5 3.1 16.8
Ens(4,3) 91.3 99.8 52.1 51.3 73.3 66.1 87.8 60.6 72.8 2.4 13.6
densenet SEA(4,3) 93.3 100.0 57.5 57.4 78.9 71.8 90.3 63.7 76.6 2.4 13.8
Ens(4,4) 84.8 94.6 55.0 54.8 76.8 73.3 87.0 57.5 73.0 6.4 21.9
Ens(4,3) 81.8 93.5 50.5 50.5 73.1 67.4 83.5 55.2 69.4 5.2 20.3
dpn SEA(4,3) 85.1 95.0 53.1 56.2 78.6 72.9 87.9 58.2 73.4 5.0 20.2
Ens(4,4) 98.1 98.4 59.2 60.4 81.0 75.8 91.7 63.2 78.5 4.5 17.9
Ens(4,3) 94.9 97.6 51.9 53.1 75.6 67.8 86.1 58.5 73.2 3.5 16.0
diverse SEA(4,3) 97.9 98.3 58.0 61.0 80.1 74.7 91.3 63.7 78.1 3.5 16.1

### v2

@@ 1 Introduction
Deep Neural Networks (DNNs) have been shown to be strikingly vulnerable to adversarial examples, i.e., maliciously crafted samples by adding small perturbations to benign inputs [Szegedy2014Intriguing, Goodfellow2015Explaining]. An intriguing property of adversarial examples that makes them threatening in the real world is their transferability [Carlini2017Towards, li2023towards, yang2025AES, Qin2022RAP, Andrew2019bugs, Zhang2022NAA], meaning that samples crafted on (known) surrogate models can also mislead (unknown) target models. Transferable attacks have been extensively studied, based on different strategies, such as gradient optimization, input transformation, and surrogate ensemble [zhao2025revisiting, gu2024survey].
@@
Among them, surrogate ensemble is widely applied due to its simplicity and easy integration with other transfer techniques [Wang2021FIA, wu2020skip, zhu2023GRA, zhu2022DRA]. In particular, there exists a well-known trade-off between attack transferability and the number of surrogate models, since using more models means lower resource efficiency regarding computational time and memory. Resource efficiency is even more critical when using transferable adversarial attacks for adversarial training-based defenses [Florian2018EnsembleAT, yang2020dverge, sitawarinpubdef].
To avoid high resource overhead, the common practice is to simply restrict the number of models to only a few, typically <5<5 [Dong2018momentum, Lin2020Nesterov, long2022Frequency]. Under such a restricted condition, although complex, optimized ensemble strategies [Xiong2021SVRE, chen2023AdaEA, tang2024ensemble, chen2024ensemble] can somehow help, the resulting transferability is still far from using many more models, according to our experimental results in Section 4.2 and a concurrent study [liu2024scaling] on scaling law of transferability. In practice, it is meaningless to restrict the number of models since various pre-trained models are easily accessible online.
In this paper, rather than proposing a new ensemble strategy, we argue that the trade-off between transferability and resource efficiency can be removed, i.e., we can adopt as many surrogate models as we want to improve transferability without sacrificing resource efficiency. This is made possible by our new interpretation of model diversity: decoupling model diversity into within-iteration and cross-iteration dimensions. In general, both dimensions account for transferability, but only within-iteration diversity (corresponding to the number of models used in each iteration) accounts for resource efficiency. Specifically, we quantify these two dimensions of diversity based on gradient similarity [kariyappa2019improving, li2020towards] to explore how they respectively contribute to the final transferability.
@@
In sum, the main contributions of this work are:
@@ 2 Related Work
Model ensemble attacks. Model ensemble attacks generally combine outputs from multiple models. Specifically, liu2017Delving averages loss from multiple models.  Dong2018momentum further proposes logits or predicted probabilities ensemble. Advanced ensemble strategies are also studied. Specifically, Xiong2021SVRE notices differences between ensemble models and proposes Stochastic Variance Reduction Ensemble (SVRE) attack to stabilize the gradient update.  chen2023AdaEA proposes an Adaptive Ensemble Attack (AdaEA), which adapts the models according to their specific contributions.  tang2024ensemble introduces a novel method of Stochastic Mini-batch black-box attack with Ensemble Reweighing (SMER) to boost the transferability.  chen2024ensemble uncovers two common weaknesses of model ensembles and relies on them to design the Common Weakness Attack (CWA). Although model ensemble attacks notably boost adversarial transferability, their practical utility is constrained by the substantial computational resources and time required for simultaneous gradient calculations across multiple models.
Gradient optimization transfer techniques. MI [Dong2018momentum] introduces momentum into adversarial attacks, stabilizing gradient update directions. NI [Lin2020Nesterov] further integrates Nesterov accelerated gradient into transfer-based attacks to better handle the history and future gradients. VT [Wang2021Variance] utilizes variance tuning to find a more stable direction for gradient updates. PGN [Ge2023flat] relied on penalizing gradient norm to achieve flat local maxima.
Input transformation transfer techniques. DI [Xie2019diversity] applies random resizing and padding operations to the input during each attack iteration to increase sample diversity and mitigate overfitting. TI [Dong2019Evading] optimizes adversarial perturbations using a set of translated images and also reduces computational costs by convolving gradients of untranslated images with a kernel matrix. SI [Lin2020Nesterov] follows a scale-invariant strategy, which optimizes the adversarial perturbations over the scale copies of the input image. SSA [long2022Frequency] transforms input images into the frequency domain and uses multiple input copies with diverse frequency features.
In this work, following the common practice in model ensemble attacks [Xiong2021SVRE, chen2024ensemble], we consider integrating the above gradient optimization and input transformation transfer techniques. Note that another type of transfer technique, model refinement [wu2020skip, zhao2025revisiting], is excluded since it focuses on model modification rather than model ensemble.
@@ 3.1 Preliminaries of Model Ensemble Attacks
[…]
where $\varepsilon$ denotes the constraint on adversarial perturbations using the $L_{\infty}$ norm. Instead of directly solving the complex optimization problem of Equation 1, the Fast Gradient Sign Method (FGSM) [Goodfellow2015Explaining] and its iterative version, I-FGSM [Kurakin2018physical], adopt the projected gradient descent.
The model ensemble is commonly used to improve adversarial transferability based on the assumption that adversarial directions learned from a diverse set of surrogate models can better generalize/transfer to unknown target models. Early studies on model ensemble attacks [liu2017Delving] propose to average the loss values of $J(f({{x}^{adv}}),y)$ in Equation 2, resulting in the ensemble loss function as:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{k}{\frac{1}{k}J({{f}_{i}}({{x}^{adv}}),y)},$ (2)
where $k$ denotes the number of ensemble models. Differently, logit output values can be averaged:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{k}{J(\frac{1}{k}{{f}_{i}}({{x}^{adv}}),y)}.$ (3)
In this paper, we adopt the logit ensemble variant due to its simplicity and effectiveness [Dong2018momentum].
@@ 3.2 Decoupling Within-iteration and Cross-iteration Model Diversity
Model ensemble attacks are known to yield improved transferability but require significant time and memory consumption due to the gradient computations on multiple models. Consequently, existing work has commonly restricted the number of models (typically <5<5) for an acceptable resource efficiency [Ge2023flat, chen2023AdaEA]. To fully investigate the impact of the number of models on the attack transferability, we conduct exploratory experiments by using 20 different models (see detailed descriptions in Section 4.1). As can be seen from the experimental results in Figure 2, using more models consistently improves the transferability, despite the performance saturation when too many models are adopted. This finding is supported by [hao2023TSEA, liu2024scaling, yao2025understanding]. Specifically, as demonstrated in [hao2023TSEA], increasing the number of models in adversarial attacks can reduce the upper bound of generalization error in empirical risk minimization.
@@
Figure 2: Attack success rates (%) of MI transferring from an ensemble of $k$ models to eight different target models. The result for TI in the Appendix D shows similar patterns.
Definition 1 Cross-iteration and Within-iteration Model Diversity. Given an iterative attack, the model diversity achieved by the number of distinct models with different architectures used throughout the entire process, $n$, is termed cross-iteration model diversity, denoted by $D_{c}$. The model diversity achieved by the number of distinct models used per iteration, $m$, is termed within-iteration model diversity, denoted by $D_{i}$. Here $m$ remains a constant in each iteration and $m\leq n$.
Quantifying diversity. To explore the relation between the two dimensions of diversity in Definition 1 and actual transferability, we further quantify the diversity. Inspired by existing work on transferable attacks/defenses, we quantify them based on gradient similarity [kariyappa2019improving, li2020towards]. Intuitively, different models yield different gradient directions on the input image. In an ensemble attack, collective gradient directions from different models enable adversarial examples to better traverse the decision boundaries of multiple models, thereby improving the possibility of transferring them to unknown target models. For within-iteration diversity, we compute the cosine similarity $S$ between gradients from different models in the same iteration:
$D_{i}=\frac{2}{m(m-1)}\sum_{1\leq i<j\leq m}S(\nabla_{x}J(f_{i}(x)),\nabla_{x}J(f_{j}(x))).$ (4)
@@
$D_{c}=S(\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t}}))}{m}},\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t+1}}))}{m}}).$ (5)
@@ 3.3 Selective Ensemble Attack (SEA)
Specifically, given $s$ easily accessible pre-trained models, denoted as $\mathcal{F}_{s}$ = $f_{1}$, $f_{2}$, $f_{i}$, …, $f_{s}$}, SEA($s$,$m$) dynamically selects $m$ models in each iteration according to the required resource constraint. A larger $s$ generally leads to a larger number of cross-iteration models $n$, but $n$ may be smaller than $s$ when the iteration number is constrained (see detailed analyses of the expectation of $n$ in Appendix A). In contrast, the conventional method Ens($s$,$m$) has to limit $n$ to $m$. For model selection, we adopt the simple random selection and compare different sampling strategies in Table 8 of Section 4.3. The implementation details of SEA are presented in Algorithm 1.
Algorithm 1 Selective Ensemble Attack (SEA)
@@
[…]
2:for $t=0$ to T1T-1 do
3:  Randomly select $m$ models from set $\mathcal{F}$${\mathcal{F}}_{m}$ = $\{f_{1},f_{2},...,f_{m}\}$ = $random.sample(\mathcal{F}_{s},m)$;
4:  Input $x$ and output ${{l}_{i}}(x_{t}^{adv})$ for i=1,2,,mi=1,2,...,m;
@@ 4.1 Experimental Setups
Dataset and models. The experiments are conducted on the ImageNet-compatible dataset [Russakovsky2015imagenet], which is widely used for evaluating adversarial examples [Dong2018momentum, Lin2020Nesterov]. This dataset consists of 1000 images from 1000 ImageNet categories (one image per category), which can be correctly classified by all the adopted models described as follows.
@@
Table 3: Attack success rates (%) of Ens vs. SEA with eight different transfer baselines. The upper bound using all 20 accessible models in each iteration is denoted as Ens(20,20), which costs $\times 5$ more computational time and memory than Ens(20,4)/SEA(20,4).
@@
For the target model, we consider 2 CNNs, 2 ViTs, and 4 Defenses. Specifically, 2 CNNs are Inception-v4 (Incv4) [Szegedy2016inc-v3] and Resnet-v2-152 (Res152) [He2016residual]. 2 ViTs are ViT-Base (ViT-B) [dosovitskiy2020image] and Swin-Base (Swin-B) [liu2021swin]. 4 Defenses are ens4-adv-Inception-v3 (ATens4), ens-adv-Inception-ResNet-v2 (ATens) [Florian2018EnsembleAT], HGD [liao2018HGD], and RS [cohen2019RS]. For the surrogate models, we select them from 20 pre-trained PyTorch models with 8 diverse architectures: resnet34/50/101/152 [He2016residual], vgg11/13/16/19 [Simonyan2015vgg], densenet121/161/169/201 [huang2017densely], fbresnet152 [He2016residual], dpn98/107/131 [Chen2017dpn], nasnetamobile [zoph2018lnasnet], senet154, and se_resnet101/152 [hu2018squeeze].
Transfer baselines and hyper-parameter settings. We integrate conventional model ensemble attacks (Ens) [Dong2018momentum] or our SEA with eight popular transfer techniques that are based on gradient optimization (i.e., MI [Dong2018momentum], NI [Lin2020Nesterov], VT [Wang2021Variance], and PGN [Ge2023flat]) or input transformation (i.e., DI [Xie2019diversity], TI [Dong2019Evading], SI [Lin2020Nesterov], and SSA [long2022Frequency]). We also consider more complex ensemble strategies, such as SVRE [Xiong2021SVRE], AdaEA [chen2023AdaEA], SMER [tang2024ensemble], and CWA [chen2024ensemble]. All the above transfer baselines have been introduced in Section 2.
@@
For clarity, we denote SEA and Ens under the ($s$,$m$) scenario as SEA($s$,$m$) and Ens($s$,$m$), respectively, where $s$ is the number of all easily accessible models and $m$ is the number of within-iteration models. When disregarding the architectural differences of models, the resource efficiency is directly proportional to $m$. For example, SEA(4,3) and Ens(4,3) are regarded as costing the same resource, and SEA(4,3) costs $1.5\times$ resource of SEA(4,2). We also report the actual time and memory costs in Tables 4, 6, 7, 8 and in Appendix D, E, F. Additionally, the analysis of model loading time for SEA and Ens is reported in Appendix C.
@@
Baseline Attack CNNs ViTs Defenses Avg.
Time
Memory
@@
Table 5: Attack success rates (%) of Ens vs. SEA on two commercial Vision APIs and large vision-language models with 100 randomly selected images.
Attack Vision APIs LVLMs
Google Baidu Qwen ChatGLM
Ens(20,4) 44 52 40 39
SEA(20,4) 58 67 52 50
Attack CNNs ViTs Defenses Avg.
Time
Memory
Incv4 Res152 ViT-B Swin-B ATens4 ATens HGD RS
Ens(20,4) 55.4 55.0 10.4 1.9 36.1 35.7 52.4 0.7 31.0 41.6 17.5
SEA(20,4) 96.3 94.5 37.6 15.0 85.4 85.6 89.9 6.7 63.9 40.7 17.9
SEA(20,3) 95.7 93.9 36.0 14.7 84.6 84.6 88.7 6.0 63.0 30.6 15.6
@@
Table 7: Transferability and resource efficiency of Ens vs. SEA with similar or diverse surrogates. In the similar cases, model ensemble is resnet18/34/50/101, vgg11/13/16/19, densenet121/161/169/201, or dpn68/98/107/131. In the diverse case, model ensemble is resnet101/vgg11/densenet161/dpn131. MI is the transfer baseline. Results for TI in the Appendix D show similar patterns.
Baseline Attack CNNs ViTs Defenses Avg.
Time
Memory
Incv4 Res152 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 85.7 91.6 51.2 51.3 74.5 63.5 90.9 62.0 71.3 1.8 7.3
Ens(4,3) 80.7 86.9 46.4 42.6 68.6 55.3 84.7 58.1 65.4 1.5 6.9
resnet SEA(4,3) 86.0 91.7 51.3 49.2 74.6 63.0 90.5 61.3 71.0 1.5 6.9
Ens(4,4) 90.1 81.0 45.4 43.8 64.2 50.9 74.5 52.3 62.8 2.6 10.6
Ens(4,3) 86.3 76.6 40.9 38.6 57.7 43.9 67.2 50.4 57.7 2.0 10.1
vgg SEA(4,3) 89.1 80.9 44.6 42.2 63.3 48.9 74.8 52.1 62.0 2.1 9.8
Ens(4,4) 91.3 93.2 58.3 56.3 78.6 71.0 91.7 63.2 75.5 3.1 16.8
Ens(4,3) 87.6 90.0 52.1 51.3 73.3 66.1 87.8 60.6 71.1 2.4 13.6
densenet SEA(4,3) 91.1 93.3 57.5 57.4 78.9 71.8 90.3 63.7 75.5 2.4 13.8
Ens(4,4) 86.9 90.2 55.0 54.8 76.8 73.3 87.0 57.5 72.7 6.4 21.9
Ens(4,3) 84.6 87.2 50.5 50.5 73.1 67.4 83.5 55.2 69.0 5.2 20.3
dpn SEA(4,3) 87.5 90.0 53.1 56.2 78.6 72.9 87.9 58.2 73.1 5.0 20.2
Ens(4,4) 91.7 93.6 59.2 60.4 81.0 75.8 91.7 63.2 77.1 4.5 17.9
Ens(4,3) 88.5 90.2 51.9 53.1 75.6 67.8 86.1 58.5 71.5 3.5 16.0
diverse SEA(4,3) 92.0 92.9 58.0 61.0 80.1 74.7 91.3 63.7 76.7 3.5 16.1